OpenStudy (anonymous):

Find the slope of the tangent line to the graph of f at the given point. f(x) = x^2 + 5x at (4, 36) A. 13 B. 21 C. 9 D. 3

OpenStudy (anonymous):

i am going to hazard a guess that you do not yet know the simple rules for finding a derivative, and have to do this by hand

OpenStudy (anonymous):

differentiate it f'(x)=2x+5 now use x=4 f'(x)=2(4)+5 f'(x)=8+5y=13 f'(x)=m=13

OpenStudy (anonymous):

let me know if i am right if you know how to find derivatives quickly, you can just about do it in your head if not, there is a bunch of work to do

OpenStudy (anonymous):

no i havent really grasped the concept yet and @saadi its x^2 not 2x

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

i assume then finding the derivative from the first principle?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you want \[\lim_{x\to 4}\frac{x^2+5x-36}{x-4}\]

OpenStudy (anonymous):

you cannot of course replace \(x\) by 4 at this step, because you will get \(\frac{0}{0}\) so the gimmick is to factor, cancel, and then replace \(x\) by 4

OpenStudy (anonymous):

and you know it MUST factor in the numerator, because if you replace \(x\) by 4 you get 0, therefore the numerator must factor as \((x-4)(\text{something})\)

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

 what happens next?